# Deep Learning for Pathological Speech: A Survey

### v1

@@ VI Intelligibility Enhancement
In pathological speech, enhancement refers to the process of improving the quality and intelligibility of speech samples affected by various impairments. The benefits of pathological speech enhancement are manifold. Firstly, the improved quality of pathological speech can be integrated into ASR systems, potentially enhancing the recognition accuracy of such speech, thereby increasing the usability of these systems for individuals with speech impairments. Secondly, these approaches can serve as data augmentation methods (r.f. VIII) for multiple applications such as pathological speech detection and automatic speech recognition.
Progress in pathological speech enhancement is limited, however, several attempts have been made by various researchers as described in the following.  [24] restructured the trajectory of formants to more closely align with the desired sound targets, thereby enhancing the accuracy of the acoustic representation. This refined formant trajectory was subsequently utilized to generate more intelligible speech, significantly improving the clarity and quality of the unintelligible pathological speech.  [153] utilized a technique where the phase of the source signal was varied across six microphone arrays to synthesize pathological speech utterances. This approach effectively increased the data samples, which were later used for data augmentation in dysarthric ASR. [25] implemented a pause insertion model along with a severity level coefficient in a neural multi-task text-to-speech system, effectively replicating and synthesizing dysarthric speech. The synthesized speech was then utilized as data augmentation for improving dysarthric speech recognition.  [154] enhanced the intelligibility of dysarthric speech by applying a methodology that includes pronunciation correction, insertion of missing sounds, devoicing of unvoiced phonemes, tempo and spectral adjustment, and removal of unwanted disfluencies. [155] demonstrated that modifying the short-term spectral content on a frame-by-frame basis can significantly improve the word-level intelligibility for patients with dysarthria. However, despite the substantial intelligibility improvements, the method’s speaker dependency poses challenges for generalizing across different speakers. In a similar speaker-dependent study, [156] employed a Kepstrum-based approach. [157] utilized multiple pronunciation patterns along with genetic algorithms and hidden Markov models [26, 27] to enhance dysarthric speech.  [158] implemented a two-stage framework in which dysarthric speech was initially recognized using an hidden Markov-based network. Subsequently, a synthesis system adapted to the dysarthric speaker was employed to produce a new speech output by correcting the mispronounced words, thereby enhancing the intelligibility of the dysarthric speech.  [159] used dynamic time warping with CNNs to improve the speech intelligibility.  [160, 161] utilized linear predictive coding mapping in conjunction with dynamic time warping between dysarthric speech and neurotypical speech to enhance dysarthric speech intelligibility in an auditory feedback mechanism. [162] employed feature space mapping through constrained maximum likelihood linear regression, aligning acoustic features from dysarthric speakers to those of neurotypical speakers. This approach demonstrated improved intelligibility of dysarthric speech, as evidenced by the perceptual evaluation of speech quality test. A similar study was conducted by [76], where dysarthric speech features were enhanced using time-delay neural network denoising autoencoders in a speaker-independent fashion. However, only a small subset of speakers were used in the evaluation, which may not generalize well to a larger pool of speakers. Motivated by the human auditory perception system, [163] introduced perceptually enhanced single frequency cepstral coefficients to improve the intelligibility of speech in dysarthric patients using i-vectors and a Gaussian mixture model with a universal background model.  [30] examined the spectro-temporal representation and showed it can be useful in improving the intelligibility of speech. [31] argued that prior methods for speech intelligibility enhancement primarily focus on speaker-dependent systems, and they introduced speech posterior feature sets designed to capture dysarthric attributes while ignoring speaker-specific characteristics. However, this assertion is not entirely accurate, as recent research has demonstrated that dysarthric speech is closely intertwined with speaker attributes [75].
In a different approach, [164] integrated Transformer-TTS, CycleVAE-VC, and LPCNet to synthesize intelligible and robust dysarthric speech, although the resulting speech lacked naturalness. [28] formulated pathological speech enhancement as a style transfer problem and proposed a generative adversarial network (GAN)-based method for dysarthric speech enhancement. Similarly, [29] used CycleGAN for converting dysarthric speech to a more intelligible speech. [165] introduced DiscoGAN to improve the speech intelligibility of dysarthric patients. This method converts pathological speech to neurotypical speech through feature mapping learned during DiscoGAN training with an adversarial and construction losses [165]. The approach showed promising results, though further validation with a larger and more diverse patient group is necessary to ensure generalizability across various pathological conditions. [166] explored a related approach and found that time-stretching techniques for dysarthric speech outperform existing state-of-the-art generative models. Their study concluded that using time-stretching with MaskCycleGAN achieves significantly better performance in dysarthric speech enhancement. However, they noted that despite these improvements, the method is not yet suitable for practical applications.
Dysarthric speech datasets typically have a limited vocabulary, but using a conversion model can expand the set of out-of-vocabulary words, which can then be used for data augmentation.  [167] presented the first end-to-end voice conversion system that trains dysarthric speech along with its corresponding textual form using a knowledge distillation approach. This method enables the replication of speech from dysarthric speakers with improved pronunciation and more precise articulation. To expand out-of-vocabulary word set, [78] utilized a voice conversion model to transform natural speech into synthetic dysarthric speech utterances. A detailed study of pathological voice conversion can be found in [168].
Neural encoder-decoder models have been shown to significantly outperform GAN-based systems in enhancing the intelligibility of dysarthric speech. However, their reliance on multi-step modules often results in inefficient and suboptimal perceptual quality. To address this issue, [169] proposed Unit-DSR, a self-supervised learning approach for dysarthric speech synthesis. It has been shown that the adversarial training framework is highly sensitive to training parameters, often leading to instability and model divergence[170, 171]. In contrast, diffusion models [172] show significant promise for applications in pathological speech enhancement, synthesis, and voice conversion. On a similar lines,  [173] employed the Glow-TTS diffusion model for generating dysarthric speech and used that as data augmentation for training a dysarthric ASR system.
@@ VII Intelligibility and Severity Assessment
Intelligibility.  Intelligibility of pathological speech is a critical indicator for evaluating the effectiveness of speech therapy and tracking the progression of various disorders. To reduce the burden of evaluating pathological speech intelligibility in clinical practice, automatic approaches have been proposed in the literature. Automatic pathological speech intelligibility assessment methods are typically categorized into two main approaches, i.e., blind and non-blind approaches [75]. In blind approaches, the objective is to assess the intelligibility of impaired speech without exploiting reference neurotypical speech data [174, 175, 176, 177, 178, 179, 180]. These approaches primarily focus on extracting acoustic features such as jitter, fundamental frequency, shimmer, formant frequencies, etc., that are believed to be closely correlated with speech intelligibility. These features are then used in regression models to estimate the intelligibility of pathological speech. Non-blind approaches, by contrast, rely on intelligible speech from neurotypical speakers as a basis for estimating the intelligibility of pathological speech [181, 182, 183, 184, 185, 186, 187, 188, 189, 190]. Such approaches typically use features extracted from ASR systems - which have been trained on large amounts of neurotypical speech - to train regression models to estimate the intelligibility of pathological speech. To avoid the burden of collecting and transcribing a large amount of neurotypical speech data required for such systems, [191] proposed the pathological short-time objective intelligibility measure (P-STOI) adapted from the speech enhancement domain. The P-STOI measure first calculates an utterance-dependent fully intelligible representation from a small set of neurotypical speakers. The intelligibility of the pathological utterance is then evaluated by quantifying its divergence from this reference representation in terms of the short-time spectral correlation. While advantageous, P-STOI requires recordings of the same utterance from intelligible neurotypical speakers, which may not always be available. To mitigate this issue, [192] developed a method to generate synthetic reference speech for assessing pathological speech intelligibility. Likewise, [193] introduced a method for invariant reference speaker pair selection to evaluate speech intelligibility. This approach leverages the SpeechSplit framework, using dynamic time warping to measure the divergence between latent representations of healthy and pathological speech utterances. In a different approach, [194] introduced sub-space based intelligibility measures based on the premise that dominant spectral patterns in pathological speech deviate significantly from those of intelligible speech. Although such measures result in a lower performance than measures exploiting neurotypical intelligible speech, they can be directly used in practical scenarios where such speech material in not available or easy to generate.
Severity.  Besides intelligibility assessment, severity assessment is another important research area where developing tools for this purpose could greatly assist in automatizing the tedious process of screening patients and categorizing them into different subgroups based on the severity level. The methods developed in the literature for this purpose can be categorized into two categories, i.e., traditional machine learning-based approaches using the Mahalanobis distance classifier [174], SVMs [195], GMMs [196], or decision trees [197], and deep learning-based approaches [141, 198, 199, 200, 201, 59, 202, 203, 204]. [205, 163] showed that i-vectors are effective in predicting the severity of pathological speech. Motivated by the human ear, [163] introduced perceptually enhanced single frequency cepstral coefficients for pathological speech intelligibility measure. Pathological speakers often exhibit irregular rhythm patterns in their speech. To leverage this characteristic, [206, 207] explored rhythm-based features for severity assessment, utilizing SVMs and MLP classifiers. Their study highlighted the potential of rhythm as a key indicator of pathological speech impairment severity, though challenges remain in capturing the full complexity of rhythm variations across different speakers and integrating these features into more advanced machine learning models. Furthermore, patients with various pathological conditions often display distinct facial expressions. To harness these visual features, [208] introduced the first audio-visual pathological severity classification system using convolutional neural networks. [200] introduced a novel training framework, demonstrating that incorporating weak speech supervision can offer significant advantages in pathological speech severity classification. To leverage metadata information such as age, gender, and type of pathological condition, [209] proposed a multi-head attention-based multi-task framework. This approach jointly optimizes severity, type, gender, and age classifications for pathological speech, thereby enhancing the robustness of latent features across these additional factors. Recently, there has been a shift towards leveraging self-supervised speech embeddings, such as Whisper and wav2vec2, in pathological speech severity measurement. This approach is gaining popularity due to the limited availability of pathological data and the ability to utilize unlabeled and metadata from other datasets, making it a promising direction for improving severity assessment in resource-constrained scenarios [77, 47].
@@ VIII Data Augmentation for
To mitigate overfitting in deep learning models dealing with the low resource pathological speech data, data augmentation techniques have been applied for various tasks. Existing approaches to data augmentation in pathological speech can be broadly classified into the following categories: traditional augmentation, voice conversion, text-to-speech (TTS) synthesis, and generative models. Drawing inspiration from the temporal and speed differences between pathological and neurotypical speech, [210, 211, 84] proposed a data augmentation technique that involves adjusting the temporal and speed characteristics of neurotypical speech to simulate pathological speech. Similarly, [212] introduced a data augmentation technique that adjusts the phonetic-level tempo of healthy speech to resemble atypical speech, and vice versa. Their findings demonstrated that the former approach is more effective for pathological ASR. [213] employed a strategy that involved combining diverse pathological speech data from multiple languages to increase the number of data samples. Their results indicated that this approach offers certain advantages in pathological ASR. Additionally, they demonstrated that integrating data from various sources can improve model performance by providing a richer and more varied dataset for training. Generative models like GANs and variational autoencoders are also commonly employed to generate pathological speech samples from healthy speech [28, 173]. These synthetic samples serve dual purposes: they can be used for speech enhancement, improving the clarity and quality of pathological speech, and for expanding the dataset, thereby enhancing the diversity of training data. By increasing the variety of pathological speech examples, these models help in mitigating overfitting and improving the performance of deep learning models in tasks such as pathological speech detection and pathological ASR [28, 173, 214, 215]. Dysarthric speech datasets typically have a limited vocabulary, but using a conversion model can expand the set of out-of-vocabulary words, which can then be used for data augmentation. To this end, [78] utilized a voice conversion model to transform natural speech into synthetic pathological speech utterances.  [83] proposed a method incorporating temporal and spectral masking combined with time warping within a contrastive learning framework to develop robust speech representations for pathological ASR. [216] employed a TTS framework to generate synthetic dysarthric speech samples as an augmentation for pathological ASR. Although this approach showed some advantages, it faced challenges in generalizing to unseen speakers. This difficulty arose from the limited variability in pathological speech data, which was insufficient to fully capture the diverse characteristics of pathological speech.  [107] investigated different data augmentation techniques in pathological ASR and demonstrated that GAN-based perturbation methods are more effective than other augmentation approaches. However, a comprehensive investigation of data augmentation methods in pathological speech detection has not yet been conducted. It would be particularly important to explore various data augmentation techniques on clean datasets such as PC-GITA, rather than on the noisy datasets such as TORGO, which has been the case for the previously reviewed literature.
IX Other Challenges and Research Directions in Pathological Speech Processing
This section describes the various challenges and possible research directions for pathological speech analysis.
IX-A Robustness in Pathological Speech Detection
@@
The methods developed so far have typically been designed and tested under specific environmental conditions and for a particular language. Each dataset used for pathological speech detection has its own unique variabilities in terms of both inter-speaker and intra-speaker differences. Additionally, different age groups exhibit distinct speaker attributes, adding complexity to pathological speech analysis. Language is another important factor; separate models have been independently developed for different languages. However, there is a need for a more robust model that is language-agnostic (to a possible degree given that pathological cues might be different in different languages), age-agnostic, accent-agnostic, and gender-agnostic to enhance overall effectiveness. Additionally, datasets such as Torgo are contaminated with noise. Researchers have shown that models trained on these datasets tend to learn environmental factors rather than focusing on genuine pathological cues [43]. To the best of our knowledge, the PC-GITA database is currently the only clean pathological dataset available under a research license. The development of pathological speech detection models robust to environmental distortions has been very limited, with only a few small studies addressing this area [217, 114, 113]. For example, [114] employed a test-time adaptation method to fine-tune pre-trained models on a validation set augmented with the test noise extracted from the test utterance. This method improves the robustness of state-of-the-art pathological speech detection methods, offering a promising solution to deploying such applications in realistic clinical settings. Similarly, [115] proposed an approach to resolve the noise disparity between the two groups of speakers in the TORGO database, such that models developed on this database learn pathology-discriminant cues instead of noise-discriminant ones. Besides robustness to environmental distortions, adversarial robustness of pathological speech detection models is another important topic and research direction. [115] explored the impact of acoustically imperceptible adversarial perturbations on deep learning-based pathological speech detection models. Results revealed a high vulnerability of such models to adversarial perturbations, with adversarial training ineffective in enhancing robustness.
IX-B Graph Neural Networks for Pathological Speech Detection

### v2

@@ VI Intelligibility Enhancement
Pathological speech enhancement refers to improving the intelligibility and quality of speech affected by impairments such as dysarthria. The benefits of such enhancement are twofold. First, enhanced speech can ease human-human and human-machine communication for pathological speakers, promoting their social and digital inclusion. Second, enhancement methods can be leveraged for data augmentation (cf. Section VIII), aiding in the development of robust models for pathological speech processing.
Early efforts in pathological speech enhancement focused on explicitly correcting articulatory and acoustic deficiencies. For instance, [33] improved intelligibility by restructuring formant trajectories to better match intended speech targets. Rudzicz [175] applied a range of techniques including pronunciation correction, phoneme insertion, tempo adjustment, and removal of disfluencies. Hosom et al. [176] modified short-term spectral features to enhance word-level intelligibility, though their approach was speaker-dependent. Lalitha et al. [177] proposed a speaker-specific Kepstrum-based method, while [178] introduced a two-stage framework that combined ASR with speech synthesis to correct mispronunciations.
Several mapping-based approaches aiming to transform dysarthric speech into more intelligible forms have also been explored. For example, [179] proposed a speech enhancement method where a CNN is trained to directly map dysarthric utterances to their control counterparts. At the feature level, techniques such as LPC mapping and frequency warping of LPC poles have been explored in [180] and [181]. Additionally, [111] investigated feature-level mapping using time-delay neural networks.
More recent advances leverage voice conversion models for pathological intelligibility enhancement. For instance, [37] framed the task as a style transfer problem and employed GANs to convert dysarthric to typical speech. Similarly, [38] utilized CycleGAN for dysarthric-to-typical speech conversion, and [182] applied DiscoGAN to map pathological and typical speech features at the acoustic level. Among these, [183] showed that time-stretching combined with MaskCycleGAN outperformed other GAN-based models in intelligibility enhancement, although the technique’s practicality remains limited due to computational complexity. In a more comprehensive system, [184] integrated Transformer-TTS, CycleVAE-VC, and LPCNet to generate highly intelligible dysarthric speech, albeit at the expense of naturalness. Additionally, [185] proposed an end-to-end voice conversion system using knowledge distillation for enhanced dysarthric speech synthesis. To avoid adversarial training instability in GAN-based voice conversion models, neural encoder-decoder architectures have also been investigated for improving pathological speech intelligibility. For instance, [186] introduced Unit-DSR, which converts dysarthric speech into discrete linguistic units and then reconstructs speech from these normalized units using a neural vocoder.
In summary, the evolution of pathological speech enhancement methods reflects a progression from early signal processing and mapping approaches to advanced generative models and neural architectures. While GAN-based models have significantly improved intelligibility, challenges remain in balancing enhancement quality, naturalness, and computational efficiency. Neural encoder-decoder frameworks show promise in addressing these challenges, marking an important direction for future research.
@@ VII Intelligibility and Severity Assessment
Intelligibility.  Intelligibility of pathological speech is a critical indicator for evaluating the effectiveness of speech therapy and tracking the progression of various disorders. To reduce the burden of evaluating pathological speech intelligibility in clinical practice, automatic approaches have been proposed in the literature. Automatic pathological speech intelligibility assessment methods are typically categorized into two main approaches, i.e., blind and non-blind approaches [110]. In blind approaches, the objective is to assess the intelligibility of impaired speech without exploiting reference neurotypical speech data [187, 188, 189, 190, 191, 192, 193]. These approaches primarily focus on extracting acoustic features such as jitter, fundamental frequency, shimmer, formant frequencies, etc., that are believed to be closely correlated with speech intelligibility. These features are then used in regression models to estimate the intelligibility of pathological speech. Non-blind approaches, by contrast, rely on intelligible speech from neurotypical speakers as a basis for estimating the intelligibility of pathological speech [194, 195, 196, 197, 198, 199, 200, 201, 202, 203]. Such approaches typically use features extracted from ASR systems, which have been trained on large amounts of control speech, to train regression models to estimate the intelligibility of pathological speech. To avoid the burden of collecting and transcribing a large amount of neurotypical speech data required for such systems, [204] proposed the pathological short-time objective intelligibility measure (P-STOI) adapted from the speech enhancement domain. The P-STOI measure first calculates an utterance-dependent fully intelligible representation from a small set of control speakers. The intelligibility of the pathological utterance is then evaluated by quantifying its divergence from this reference representation in terms of the short-time spectral correlation. While advantageous, P-STOI requires recordings of the same utterance from intelligible control speakers, which may not always be available. To mitigate this issue, [205] developed a method to generate synthetic reference speech for assessing pathological speech intelligibility. In a different approach, [206] introduced sub-space based intelligibility measures based on the premise that dominant spectral patterns in pathological speech deviate significantly from those of intelligible speech. Although such measures result in a lower performance than measures exploiting neurotypical intelligible speech, they can be directly used in practical scenarios where such speech material in not available or easy to generate.
Severity.  Besides intelligibility assessment, severity assessment is another important research area where developing tools for this purpose could greatly assist in automatizing the tedious process of screening patients and categorizing them into different subgroups based on the severity level. The methods developed in the literature for this purpose can be categorized into two categories, i.e., traditional machine learning-based approaches using the Mahalanobis distance classifier [187], SVMs [207], GMMs [208], or decision trees [209], and deep learning-based approaches [210, 211, 212, 90, 213, 214, 215]. While deep learning approaches aim to automatically extract acoustic cues correlated with severity from raw or minimally processed speech signals, traditional machine learning approaches for severity assessment rely on (clinically informed) handcrafted acoustic features. For example, motivated by auditory processing knowledge, [216] introduced perceptually enhanced single frequency cepstral coefficients for assessing the severity of pathological speech. Vásquez-Correa et al. [217] showed that articulation features extracted from continuous speech signals to create i-vectors were advantageous in quantifing the dysarthria severity level. Based on the knowledge that pathological speakers often exhibit irregular rhythm patterns in their speech, [218, 219] explored rhythm-based features for severity assessment.
Besides speech impairments, patients with various pathological conditions often display distinct facial expressions. To harness these visual features, [220] introduced the first audio-visual pathological severity classification system using CNNs. To leverage metadata information such as age, sex, and type of pathological condition, [221] proposed a multi-head attention-based MTL framework. This approach jointly optimizes severity, type, sex, and age classifications for pathological speech, thereby enhancing the robustness of latent features across these additional factors. Recently, there has been growing interest in using SSL embeddings to measure pathological speech severity. This approach is promising due to the scarcity of labeled pathological data and the ability to leverage unlabeled data and metadata from other datasets, making it well-suited for resource-constrained settings [112, 79].
@@ VIII Data Augmentation for
To mitigate overfitting in deep learning models dealing with the low resource pathological speech data, data augmentation techniques have been used for various tasks. Existing approaches to data augmentation in pathological speech are broadly based on traditional strategies (such as incorporating noise, reverberation, or multiple datasets), perturbation strategies, voice conversion, and text-to-speech (TTS) synthesis.
@@
Takashima et al. [222] employed a strategy that involved combining diverse pathological speech data from multiple languages to increase the number of data samples. Their results indicate that such a traditional data augmentation approach can be advantageous for pathological ASR. Based on the temporal and speed differences between pathological and control speech, [223, 224, 119] have explored vocal tract length perturbation, tempo perturbation, and speed perturbation as data augmentation approaches for pathological ASR. Similarly, [225] introduced a data augmentation technique that adjusts the phonetic-level tempo of healthy speech to resemble atypical speech, and vice versa. Their findings demonstrated that the former approach is more effective for pathological ASR. More recently, voice conversion and TTS systems are also commonly employed to generate pathological speech from healthy speech [37, 226, 227]. These synthetic samples serve a dual purpose, i.e., they can be used for speech enhancement, improving the clarity and quality of pathological speech; and they can be used for expanding the dataset, thereby enhancing the diversity of training data. Increased data diversity is beneficial to avoid overfitting and improve the performance of deep learning models in tasks such as pathological speech detection and pathological ASR [37, 226, 228, 229]. Since pathological speech datasets typically have a limited vocabulary, using a voice conversion or TTS model can also be used to expand the set of out-of-vocabulary words [113]. Recently, [147] investigated different data augmentation strategies in pathological ASR, demonstrating that GAN-based conversion methods are more effective than perturbation-based augmentation approaches. However, a comprehensive investigation of the advantages of all data augmentation strategies in various pathological speech processing tasks using multiple datasets is still lacking.
IX Challenges and Research Directions in Pathological Speech Processing
In clinical settings, the integration of automated speech processing systems for pathological speech analysis is crucial for advancing the diagnosis, therapy, and monitoring of various disorders. While standardized clinical scales are invaluable in perceptually evaluating pathological speech, the incorporation of automated systems can offer additional benefits in terms of objectivity, efficiency, and real-time feedback. Tools such as VoxTester [230], for instance, provide clinicians with quantifiable speech metrics, including articulation precision and speech rate, which help monitor disease progression and assess the effectiveness of interventions. However, to fully realize the potential of these tools, further research is needed to ensure that they can seamlessly integrate into clinical workflows, offering clinicians user-friendly, reliable, and interpretable outputs. Addressing issues such as clinician training, data security, patient consent, and adaptability to various clinical settings is crucial for the real-world adoption of these systems. Moreover, creating systems that can operate longitudinally, i.e., tracking speech performance over time to capture subtle changes in speech function, would significantly enhance therapeutic outcomes. We believe that addressing the challenges and research directions outlined in the remainder of this section will be key to enabling the seamless integration of these technologies into clinical practice in the future.
IX-A Impact of Speech Mode
The large majority of previously reviewed deep learning-based pathological speech approaches have been proposed and validated on controlled speech tasks. Controlled speech, also known as non-spontaneous speech, involves utterances produced within a structured context, typically requiring participants to repeat phonetically balanced, carefully crafted texts. This mode is designed to elicit specific pathological biomarkers by standardizing the motor planning and execution demands. Tasks may include reading aloud or repeating scripted phrases. In contrast, spontaneous speech consists of unplanned utterances, such as storytelling or casual conversation, which reflect real-world communicative behavior. It places greater demands on cognitive planning, articulation, and natural language generation, and may therefore reveal more authentic or varied pathological cues. While this distinction has received some attention in the context of detection, like in [92], its implications extend across other subfields such as severity assessment, intelligibility prediction, enhancement, and recognition. For instance, severity assessment models trained only on controlled speech may generalize poorly to real-world settings. Similarly, enhancement models may be tuned to the acoustic patterns of controlled tasks, failing to capture the variability in spontaneous speech. Incorporating spontaneous speech more broadly into pathological speech research could enhance practical utility and performance across these subfields. Given its ease of collection and alignment with natural communication, spontaneous speech provides a more suitable and informative basis for training and evaluating models [92, 231, 232]. Future work should consider systematically analyzing the effect of speech mode across multiple application areas to ensure generalizability and real-world effectiveness.
IX-B Robustness in Pathological Speech Detection
@@
The methods developed so far for automatic pathological speech detection have typically been designed and tested under specific environmental conditions and for a particular language. Each dataset used for pathological speech detection has its own unique variabilities in terms of both inter-speaker and intra-speaker differences. Additionally, different age groups exhibit distinct speaker attributes, adding complexity to pathological speech analysis. Language is another important factor; separate models have been independently developed for different languages. However, there is a need for a more robust model that is language-agnostic (to a possible degree given that pathological cues might be different in different languages), age-agnostic, accent-agnostic, and sex-agnostic to enhance overall effectiveness [151]. Additionally, datasets such as TORGO are contaminated with noise. Researchers have shown that models trained on these datasets tend to learn environmental factors rather than focusing on genuine pathological cues [70]. The development of pathological speech detection models robust to environmental distortions has been very limited, with only a few small studies addressing this area [233, 234, 71, 235]. For example, [234] employed a test-time adaptation method to fine-tune pre-trained models on a validation set augmented with the test noise extracted from the test utterance. This method improves the robustness of state-of-the-art pathological speech detection methods, offering a promising solution to deploying such applications in realistic clinical settings. Similarly, [236] proposed an approach to resolve the noise disparity between the two groups of speakers in the TORGO database, such that models developed on this database learn pathology-discriminant cues instead of noise-discriminant ones. Besides robustness to environmental distortions, adversarial robustness of pathological speech detection models is another important topic and research direction. The impact of acoustically imperceptible adversarial perturbations on deep learning-based pathological speech detection models has been explored in [71]. Results revealed a high vulnerability of such models to adversarial perturbations, with adversarial training ineffective in enhancing robustness.
IX-C Improving the Performance of Automatic Pathological Speech Analysis